OpenStudy (anonymous):

How would I soulve 6x+10y=12 -3x+y=6 using elimination

OpenStudy (btaylor):

multiply both sides of the bottom equation by 2 to get: 6x+10y=-12 -6x+2y = 12 now add straight down, and you get 12y=0. So y=0. Now plug 0 in for y in either equation to get x.

OpenStudy (anonymous):

x=-2?

OpenStudy (btaylor):

no, x=2

OpenStudy (anonymous):

ohhh. Is that all I do, is the problem done?

OpenStudy (btaylor):

yes.